# T versus CP effects in DUNE and T2HK

### v1

@@
August 6, 2025
@@ 1 Introduction
A genuine property of weak interactions with three fermion generations is the violation of the time reversal (T) or equivalently, the charge-parity (CP) conjugation symmetries of the fundamental theory [1]. This phenomenon may also take place in neutrino oscillations [2, 3, 4], provided suitable non-trivial complex phases are present in the theory. Within the standard three-flavour paradigm, fundamental T and CP violation are equivalent and both are controlled by a single complex in the PMNS [5, 6] mixing matrix, conventionally denoted as $\delta_{\rm CP}$. Current data on neutrino oscillations [7, 8, 9, 10] show first indications for preferred regions for $\delta_{\rm CP}$ and the search for effects of such complex phases is the main science goal of the next generation long-baseline oscillation experiments T2HK [11] and DUNE [12, 13].
@@
Traditionally, the T transformation in neutrino oscillations is related to the exchange of initial and final neutrino flavours: $\mathcal{T}[P_{\alpha\to\beta}(L)]=P_{\beta\to\alpha}(L)$ see e.g., refs. [2, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] for an incomplete list of references. Due to the experimental difficulty to produce an electron or tau neutrino beam, this method cannot be pursued with conventional pion-based neutrino beams and probably requires the advent of a muon-based neutrino factory, see e.g. [25] for a recent paper. Here, instead, we follow the approach of [28, 29], which is based on the well-known observation that the T transformation is equivalent to the formal replacement of the baseline $L\to-L$ but keeping the oscillation channel unchanged, i.e., $\mathcal{T}[P_{\alpha\to\beta}(L)]=P_{\alpha\to\beta}(-L)$ (see also the appendix of ref. [29] for a discussion of this property). Hence, we can search for T violation by looking for an $L$-odd component of the transition probability $P_{\nu_{\mu}\to\nu_{e}}(L)$, considered as a function of $L$ but at fixed neutrino energy.
@@
In Refs. [28, 29] the emphasis has been put on model-independent measures for T violation. In the present paper we relax these ambitions with regard to model-independence and apply the approach of [28] to the case of the standard unitary three-flavour mixing model. As we will show, this significantly increases the sensitivity, when the present knowledge [9, 10] about the mass and mixing parameters is taken into account. Technically, the analysis becomes very similar to the standard $\delta_{\rm CP}$ determination. However, we argue that when results are interpreted in terms of the appearance probability as a function of baseline, we see explicitly the effect of T-odd contributions to the probability. We will study the potential of DUNE and T2HK to explore this effect, and contrast it to “CP observables” related to the comparison of neutrino versus anti-neutrino transition probabilities. As we will show, even with an appropriately chosen single energy bin with only neutrino beam running, DUNE has significant sensitivity to T violation. A crucial role in this respect will come from information from the second oscillation maximum, emphasizing the importance of the low energy part of the DUNE event spectrum, see e.g. [30, 31, 32, 33, 34, 35, 36, 37] for previous studies on the 2nd oscillation maximum. In this sense, DUNE has better sensitivity to T violation (based on data only from neutrino beam mode), whereas T2HK preferably determines CP violation (based on the comparison of neutrino and anti-neutrino data).
@@
The outline of the paper is as follows. In section 2 we briefly introduce the oscillation formalism and give technical details on our numerical simulation of the DUNE and T2HK experiments. In section 3 we discuss the numerical results about T violation from $L$ dependence of the oscillation probability: we identify the most sensitive energy range, emphasizing the importance of the 2nd oscillation maximum in DUNE as well as the synergy with T2HK, and study the dependence of the sensitivity to prior knowledge on oscillation parameters as well as on the exposure. In section 4 we discuss the sensitivity to T violation compared to CP violation (based on the neutrino/anti-neutrino comparison) and work out the complementarity between the DUNE and T2HK experiments in this respect. In section 5 we consider the model-independent observable $X_{T}$ introduced in ref. [29], which is given by the difference of the oscillation probability at the baselines of the DUNE and T2HK experiments, evaluated at the same neutrino energy: $X_{T}\equiv P_{\nu_{\mu}\to\nu_{e}}(L_{2})-P_{\nu_{\mu}\to\nu_{e}}(L_{1})$. Under certain conditions, a negative value of $X_{T}$ is a model-independent signal of T violation. Here we will investigate the sensitivity of DUNE and T2HK to the $X_{T}$ test within the standard three-flavour oscillation framework. We conclude in section 6. In appendix A we offer further details on the energy dependence of T and CP violation by studying the bin-per-bin sensitivity of the two experiments.
@@ 2.1 Transition probabilities
Let us briefly review the theoretical framework and fix our notation. We focus on the $\nu_{\mu}\to\nu_{e}$ appearance probability $P_{\nu_{\mu}\to\nu_{e}}$, with the $L$ and $E_{\nu}$ dependence often left implicit. We approximate the matter density along the neutrino path being constant and the same for T2HK and DUNE. The impact of deviations from this approximation on the T-violation test has been quantified in [38] and found to be small, see also [26]. Then the oscillation probability is obtained as
@@
where $\lambda_{i}$ are the eigenvalues of the effective Hamiltonian including the matter potential [39], and $U_{\alpha i}^{m}$ is the effective mixing matrix in matter. Because $U_{\alpha i}^{m}$ is unitary, only two out of the three $c_{i}$ coefficients are independent, which we take as $c_{2},c_{3}$. One can split the probability into T-even and T-odd parts:
@@
$\phi_{ij}\equiv\frac{\lambda_{i}-\lambda_{j}}{2}L=\frac{\Delta m^{2}_{ij,{\rm eff}}(E_{\nu})L}{2E_{\nu}}\,$, (2.5)
with $\Delta m^{2}_{ij,{\rm eff}}(E_{\nu})$ denoting the effective mass-squared differences in matter. Note that in eqs. 2.3 and 2.4 we see explicitly the even and odd behaviour under $L\to-L$, respectively.
Considering now the transition probability as a function of baseline for fixed $E_{\nu}$, T violation can be established if data cannot be fitted with $P_{\rm even}(L)$ alone. In Ref. [28], this test has been formulated in a general way, imposing no prior knowledge on the coefficients $c_{i}$ in order to remain as model-independent as possible. In contrast, in the following we will specialize to the standard three-flavour scenario by using the specific expressions for $c_{i}$ and $\phi_{ij}$ implied by eqs. 2.1 and 2.5, and impose the prior knowledge on the three mixing angles $\theta_{ij}$, mass-squared differences $\Delta m^{2}_{ij}$, as well as the Standard Model matter potential. In this way, the search for T violation can be applied even with a single measurement at suitable chosen baseline and energy.
@@ 2.2 Analysis details
DUNE: The Deep Underground Neutrino Experiment (DUNE) will generate an intense beam of neutrinos at Fermi National Accelerator Laboratory (Fermilab) in Illinois and direct it toward the far detector located approximately 1300 km away at the Sanford Underground Research Facility (SURF) in South Dakota. We have adopted the experimental setup described in the Technical Design Report (TDR) [12, 40], including systematic uncertainties and detector efficiencies. The TDR configuration consists of a far detector made of a 40 kton Liquid Argon Time Projection Chamber, offering exceptional tracking and particle identification capabilities. The current plan follows a staged deployment with up to four 17 kton detector modules. DUNE will use a high-intensity 120 GeV proton beam with a power of 1.2 MW, expected to deliver $1.1\times 10^{21}$ protons on target (P.O.T) per year. For our analysis, the low-energy tail of the event spectrum in DUNE is relevant, which suffers from low event numbers, see [29] for a discussion. Therefore, results are largely limited by DUNE statistical uncertainties. As our default exposure we assume 336 kt MW yr, which should be achieved roughly after 7 years of operation [41]. As we focus on the sensitivity for neutrinos, we assume that the full exposure is taken in the neutrino beam mode. Anti-neutrino running is considered in section 4 and appendix A. For the energy resolution we assume an improved performance, as proposed in Refs. [36, 42] (see also [43, 44]). We consider a Gaussian detector resolution with $\sigma=\alpha E_{\nu}+\beta\sqrt{E_{\nu}}+\gamma$, where $E_{\nu}$ is the neutrino energy in GeV. Our default assumption is $(\alpha,\beta,\gamma)=(0.045,0.001,0.048)$, with units in GeV. This choice corresponds to the resolution from [42, 44], see also [29] for a discussion.
@@
T2HK: The Tokai-to-hyper-Kamiokande (T2HK) experiment is a next-generation off-axis accelerator-based neutrino project designed to study neutrino oscillations with a 295 km baseline. It uses the existing 30 GeV proton beam from the J-PARC accelerator facility in Tokai, Japan, to produce intense beams of neutrinos and antineutrinos. The experiment will employ a massive Water Cherenkov far detector located at the Hyper-Kamiokande (HK) site, with a fiducial volume of 187 kilotons. To assess the physics reach of T2HK, we follow the experimental specifications detailed in the Hyper-Kamiokande design report [11]. The total integrated exposure is expected to be $1.3~\rm{MW}\times 10\times 10^{7}$ seconds, corresponding to $2.7\times 10^{22}$ protons on the target (PT). Our default exposure for T2HK it is 608 kt MW yr, corresponding roughly to 2.5 yr of running in the neutrino mode. As the relevant energy window for the T-violation test is located close to the maximum of the event spectrum, T2HK statistical errors are subdominant. In our analysis, we adopt a simplified treatment of systematic uncertainties, including an uncorrelated 5% (3.5%) signal normalization error, a 10% background normalization error, and a 5% energy calibration error. For the energy resolution we use the same parameterization as for DUNE as mentioned above, with the values $(\alpha,\beta,\gamma)=(0.12,0.07,0.0)$, which match the resolution provided in the design report [11].
@@
In order to estimate the sensitivity to T-violation in neutrino oscillations, we use the GLoBES software package [45, 46], implementing necessary modifications to accommodate our specific analysis framework. Our study focuses exclusively on the appearance channel, as this channel provides the most direct probe of T-asymmetries in the neutrino sector. In our analysis, we treat the mass-squared splittings $\Delta m^{2}_{21}$ and $\Delta m^{2}_{31}$ as fixed parameters, constrained by high-precision external measurements, such as the disappearance channel from DUNE/T2HK or the JUNO reactor experiment [47]. The neutrino mass ordering is assumed to be known and fixed to normal ordering. Consequently, the only free parameters in our oscillation fit are the three mixing angles $\theta_{12}$, $\theta_{13}$ and $\theta_{23}$.
@@
For the T-violation analysis we only use simulated data in the neutrino beam mode. In principle, an analogous analysis can be performed also with anti-neutrinos. However, due to the sizeable intrinsic neutrino component of the anti-neutrino beams both for T2HK and DUNE, it is not possible to extract the anti-neutrino transition probability in a clean way from the data. Hence, data in the anti-neutrino beam mode does not allow a straight-forward interpretation in terms of $L$-odd versus $L$-even components and T-violating observables are significantly diluted [29]. Therefore, we restrict the T violation analysis to the neutrino mode only, where the intrinsic anti-neutrino contamination is sufficiently small to be neglected.
@@
Table 1: Best-fit values and $1\sigma$ uncertainties of the standard three-flavour neutrino parameters [9] adopted in the numerical analysis.

### v2

@@
December 10, 2025
Time reversal (T) symmetry violations in neutrino oscillations imply the presence of an $L$-odd component in the transition probability at fixed neutrino energy, with $L$ denoting the distance between neutrino source and detector. Within the standard three-flavour framework, we show that the combination of the transition probabilities determined at the DUNE and T2HK experiments can establish the presence of an $L$-odd component, and therefore provide sensitivity to T violation, up to $4\sigma$ significance. The optimal neutrino energy window is from 0.68 to 0.92 GeV, and therefore a crucial role is played by the low-energy part of the DUNE event spectrum covering the second oscillation maximum. We compare the sensitivity to T violation based on this energy range using neutrino data only with the more traditional search for charge-parity (CP) violation based on the comparison of neutrino versus anti-neutrino beam data. We show that for DUNE it is advantageous to run in neutrino mode only, i.e., searching for T violating effects, whereas T2HK is more sensitive to CP violation, comparing neutrino and anti-neutrino data. Hence, the two experiments offer complementary methods to determine the complex phase in the PMNS mixing matrix.
@@ 1 Introduction
A genuine property of weak interactions with three fermion generations is the violation of the time reversal (T) or equivalently, the charge-parity (CP) conjugation symmetries of the fundamental theory [Kobayashi:1973fv]. This phenomenon may also take place in neutrino oscillations [Cabibbo:1977nk, Bilenky:1980cx, Barger:1980jm], provided suitable non-trivial complex phases are present in the theory. Within the standard three-flavour paradigm, fundamental T and CP violation are equivalent and both are controlled by a single complex phase in the PMNS [Pontecorvo:1967fh, Maki:1962mu] mixing matrix, conventionally denoted as $\delta_{\rm CP}$. Current data on neutrino oscillations [T2K:2023smv, NOvA:2021nfi, Esteban:2024eli, Capozzi:2025wyn] show first indications for preferred regions for $\delta_{\rm CP}$ and the search for effects of such complex phases is the main science goal of the next generation long-baseline oscillation experiments T2HK [Hyper-Kamiokande:2018ofw] and DUNE [DUNE:2020jqi, Abi:2020wmh].
@@
Traditionally, the T transformation in neutrino oscillations is related to the exchange of initial and final neutrino flavours: $\mathcal{T}[P_{\alpha\to\beta}(L)]=P_{\beta\to\alpha}(L)$ see e.g., refs. [Cabibbo:1977nk, Kuo:1987km, Krastev:1988yu, Toshev:1989vz, Toshev:1991ku, Arafune:1996bt, Parke:2000hu, Akhmedov:2001kd, Schwetz:2007py, Xing:2013uxa, Petcov:2018zka, Bernabeu:2019npc, Kitano:2024kdv, Bitter:2024eka, Yasuda:2024lwg] for an incomplete list of references. Due to the experimental difficulty to produce an electron or tau neutrino beam, this method cannot be pursued with conventional pion-based neutrino beams and probably requires the advent of a muon-based neutrino factory, see e.g. [Kitano:2024kdv] for a recent paper. Here, instead, we follow the approach of [Schwetz:2021cuj, Chatterjee:2024jzt], which is based on the well-known observation that the T transformation is equivalent to the formal replacement of the baseline $L\to-L$ but keeping the oscillation channel unchanged, i.e., $\mathcal{T}[P_{\alpha\to\beta}(L)]=P_{\alpha\to\beta}(-L)$ (see also the appendix of ref. [Chatterjee:2024jzt] for a discussion of this property). Hence, we can search for T violation by looking for an $L$-odd component of the transition probability $P_{\nu_{\mu}\to\nu_{e}}(L)$, considered as a function of $L$ but at fixed neutrino energy.
@@
In Refs. [Schwetz:2021cuj, Chatterjee:2024jzt] the emphasis has been put on model-independent measures for T violation. In the present paper we relax these ambitions with regard to model-independence and apply the approach of [Schwetz:2021cuj] to the case of the standard unitary three-flavour mixing model. As we will show, this significantly increases the sensitivity, when the present knowledge [Esteban:2024eli, Capozzi:2025wyn] about the mass and mixing parameters is taken into account. Technically, the analysis becomes very similar to the standard $\delta_{\rm CP}$ determination. However, we argue that when results are interpreted in terms of the appearance probability as a function of baseline, we see explicitly the effect of T-odd contributions to the probability. We will study the potential of DUNE and T2HK to explore this effect, and contrast it to “CP observables” related to the comparison of neutrino versus anti-neutrino transition probabilities. As we will show, even with an appropriately chosen single energy bin with only neutrino beam running, DUNE has significant sensitivity to T violation. A crucial role in this respect will come from information from the second oscillation maximum, emphasizing the importance of the low energy part of the DUNE event spectrum, see e.g. [Ishitsuka:2005qi, Huber:2010dx, Coloma:2011pg, Qian:2013nhp, Wildner:2015yaa, Hyper-Kamiokande:2016srs, DeRomeri:2016qwo, Rout:2020emr] for previous studies on the 2nd oscillation maximum. In this sense, DUNE has better sensitivity to T violation (based on data only from neutrino beam mode), whereas T2HK preferably determines CP violation (based on the comparison of neutrino and anti-neutrino data).
@@
The outline of the paper is as follows. In section 2 we briefly introduce the oscillation formalism and give technical details on our numerical simulation of the DUNE and T2HK experiments. In section 3 we discuss the numerical results about T violation from $L$ dependence of the oscillation probability: we identify the most sensitive energy range, emphasizing the importance of the 2nd oscillation maximum in DUNE as well as the synergy with T2HK, and study the dependence of the sensitivity to prior knowledge on oscillation parameters as well as on the exposure. In section 4 we discuss the sensitivity to T violation compared to CP violation (based on the neutrino/anti-neutrino comparison) and work out the complementarity between the DUNE and T2HK experiments in this respect. In section 5 we consider the model-independent observable $X_{T}$ introduced in ref. [Chatterjee:2024jzt], which is given by the difference of the oscillation probability at the baselines of the DUNE and T2HK experiments, evaluated at the same neutrino energy: $X_{T}\equiv P_{\nu_{\mu}\to\nu_{e}}(L_{2})-P_{\nu_{\mu}\to\nu_{e}}(L_{1})$. Under certain conditions, a negative value of $X_{T}$ is a model-independent signal of T violation. Here we will investigate the sensitivity of DUNE and T2HK to the $X_{T}$ test within the standard three-flavour oscillation framework. We conclude in section 6. In appendix A we offer further details on the energy dependence of T and CP violation by studying the bin-per-bin sensitivity of the two experiments whereas in appendix B we justify our approximation concerning the matter density.
@@ 2.1 Transition probabilities
Let us briefly review the theoretical framework and fix our notation. We focus on the $\nu_{\mu}\to\nu_{e}$ appearance probability $P_{\nu_{\mu}\to\nu_{e}}$, with the $L$ and $E_{\nu}$ dependence often left implicit. We approximate the matter density along the neutrino path being constant and the same for T2HK and DUNE. The impact of deviations from this approximation on the T-violation test has been quantified in [Schwetz:2021thj] and found to be small, see also [Bitter:2024eka]. Then the oscillation probability is obtained as
@@
where $\lambda_{i}$ are the eigenvalues of the effective Hamiltonian including the matter potential [Wolfenstein:1977ue], and $U_{\alpha i}^{m}$ is the effective mixing matrix in matter. Because $U_{\alpha i}^{m}$ is unitary, only two out of the three $c_{i}$ coefficients are independent, which we take as $c_{2},c_{3}$. One can split the probability into T-even and T-odd parts:
@@
$\phi_{ij}\equiv\frac{\lambda_{i}-\lambda_{j}}{2}L=\frac{\Delta m^{2}_{ij,{\rm eff}}(E_{\nu})L}{4E_{\nu}}\,$, (2.5)
with $\Delta m^{2}_{ij,{\rm eff}}(E_{\nu})$ denoting the effective mass-squared differences in matter. Note that in eqs. 2.3 and 2.4 we see explicitly the even and odd behaviour under $L\to-L$, respectively. The factor ${\rm Im}[c_{2}^{*}c_{3}]={\rm Im}[U_{\mu 2}^{m}U_{e2}^{m*}U_{\mu 3}^{m*}U_{e3}^{m}]$ in eq. 2.4 corresponds to the Jarlskog invariant [Jarlskog:1985ht] for mixing in matter, see also [Freund:2001pn, Denton:2019yiw] for discussions.
Considering now the transition probability as a function of baseline for fixed $E_{\nu}$, T violation can be established if data cannot be fitted with $P_{\rm even}(L)$ alone. In Ref. [Schwetz:2021cuj], this test has been formulated in a general way, imposing no prior knowledge on the coefficients $c_{i}$ in order to remain as model-independent as possible. In contrast, in the following we will specialize to the standard three-flavour scenario by using the specific expressions for $c_{i}$ and $\phi_{ij}$ implied by eqs. 2.1 and 2.5, and impose the prior knowledge on the three mixing angles $\theta_{ij}$, mass-squared differences $\Delta m^{2}_{ij}$, as well as the Standard Model matter potential. In this way, the search for T violation can be applied even with a single measurement at suitable chosen baseline and energy.
@@ 2.2 Analysis details
DUNE: The Deep Underground Neutrino Experiment (DUNE) will generate an intense beam of neutrinos at Fermi National Accelerator Laboratory (Fermilab) in Illinois and direct it toward the far detector located approximately 1300 km away at the Sanford Underground Research Facility (SURF) in South Dakota. We have adopted the experimental setup described in the Technical Design Report (TDR) [DUNE:2020jqi, DUNE:2021cuw], including systematic uncertainties and detector efficiencies. The TDR configuration consists of a far detector made of a 40 kton Liquid Argon Time Projection Chamber, offering exceptional tracking and particle identification capabilities. The current plan follows a staged deployment with up to four 17 kton detector modules. DUNE will use a high-intensity 120 GeV proton beam with a power of 1.2 MW, expected to deliver $1.1\times 10^{21}$ protons on target (P.O.T) per year. For our analysis, the low-energy tail of the event spectrum in DUNE is relevant, which suffers from low event numbers, see [Chatterjee:2024jzt] for a discussion. Therefore, results are largely limited by DUNE statistical uncertainties. As our default exposure we assume 336 kt MW yr, which should be achieved roughly after 7 years of operation [IRN-talk]. As we focus on the sensitivity for neutrinos, we assume that the full exposure is taken in the neutrino beam mode. Anti-neutrino running is considered in section 4 and appendix A. For the energy resolution we assume an improved performance, as proposed in Refs. [DeRomeri:2016qwo, Friedland:2018vry] (see also [Kopp:2024lch, Chatterjee:2021wac]). We consider a Gaussian detector resolution with $\sigma=\alpha E_{\nu}+\beta\sqrt{E_{\nu}}+\gamma$, where $E_{\nu}$ is the neutrino energy in GeV. Our default assumption is $(\alpha,\beta,\gamma)=(0.045,0.001,0.048)$, with units in GeV. This choice corresponds to the resolution from [Friedland:2018vry, Chatterjee:2021wac], see also [Chatterjee:2024jzt] for a discussion.
@@
T2HK: The Tokai-to-hyper-Kamiokande (T2HK) experiment is a next-generation off-axis accelerator-based neutrino project designed to study neutrino oscillations with a 295 km baseline. It uses the existing 30 GeV proton beam from the J-PARC accelerator facility in Tokai, Japan, to produce intense beams of neutrinos and antineutrinos. The experiment will employ a massive Water Cherenkov far detector located at the Hyper-Kamiokande (HK) site, with a fiducial volume of 187 kilotons. To assess the physics reach of T2HK, we follow the experimental specifications detailed in the Hyper-Kamiokande design report [Hyper-Kamiokande:2018ofw]. The total integrated exposure is expected to be $1.3~\rm{MW}\times 10\times 10^{7}$ seconds, corresponding to $2.7\times 10^{22}$ protons on the target (PT). Our default exposure for T2HK it is 608 kt MW yr, corresponding roughly to 2.5 yr of running in the neutrino mode. As the relevant energy window for the T-violation test is located close to the maximum of the event spectrum, T2HK statistical errors are subdominant. In our analysis, we adopt a simplified treatment of systematic uncertainties, including an uncorrelated 5% (3.5%) signal normalization error, a 10% background normalization error, and a 5% energy calibration error. For the energy resolution we use the same parameterization as for DUNE as mentioned above, with the values $(\alpha,\beta,\gamma)=(0.12,0.07,0.0)$, which match the resolution provided in the design report [Hyper-Kamiokande:2018ofw].
@@
In order to estimate the sensitivity to T-violation in neutrino oscillations, we use the GLoBES software package [Huber:2004ka, Huber:2007ji], implementing necessary modifications to accommodate our specific analysis framework. Our study focuses exclusively on the appearance channel, as this channel provides the most direct probe of T-asymmetries in the neutrino sector. In our analysis, we treat the mass-squared splittings $\Delta m^{2}_{21}$ and $\Delta m^{2}_{31}$ as fixed parameters, constrained by high-precision external measurements, such as the disappearance channel from DUNE/T2HK or the JUNO reactor experiment [JUNO:2022mxj]. The neutrino mass ordering is assumed to be known and fixed to normal ordering. Consequently, the only free parameters in our oscillation fit are the three mixing angles $\theta_{12}$, $\theta_{13}$ and $\theta_{23}$.
@@
For the T-violation analysis we only use simulated data in the neutrino beam mode. In principle, an analogous analysis can be performed also with anti-neutrinos. However, due to the sizeable intrinsic neutrino component of the anti-neutrino beams both for T2HK and DUNE, it is not possible to extract the anti-neutrino transition probability in a clean way from the data. Hence, data in the anti-neutrino beam mode does not allow a straight-forward interpretation in terms of $L$-odd versus $L$-even components and T-violating observables are significantly diluted [Chatterjee:2024jzt]. Therefore, we restrict the T violation analysis to the neutrino mode only, where the intrinsic anti-neutrino contamination is sufficiently small to be neglected.
@@
Table 1: Best-fit values and $1\sigma$ uncertainties of the standard three-flavour neutrino parameters [Esteban:2024eli] adopted in the numerical analysis.